Tom can mow ½ the lawn in an hour when working alone. Jerry can mow ⅔ of the lawn in an hour when working alone. How long will it take each person in the scenario to mow the lawn working alone? How long would it take them to mow the lawn if they worked together?
Answer by Boreal(15235)

It takes Tom 1/(1/2) or 2 hours to do it alone.
It takes Jerry 1/(2/3) or 3/2 hours to do it alone.
Tom does (x/2) of the lawn per hour and Jerry does 2x/3 of the lawn per hour
together (x/2)+(2x/3)=1
and multiplying by 6, 3x+4x=6
7x=6
x=6/7 of an hour together.
Tom will do 3/7 of it and Jerry will do 4/7 of it.